# TBStar-Edit: From Image Editing Pattern Shifting to Consistency Enhancement

### v1

@@ 2.2 Editing Pairs Construction Methods
Constructing editing pairs from collected raw image data is the core process in building image editing datasets. Given the wide variety of image editing tasks, each with distinct characteristics and requirements, we systematically categorize these tasks and, accordingly, establish four different methods for constructing editing pairs: the Editing Expert Model Based Method, the Template Based Method, the In-Context Generation Based Method, and the LoRA Based Method.
@@ 2.2.3 In-Context Generation Based Method
2.2.4 LoRA Hu et al. (2022) Based Method
@@ 3.1 Model Architecture
Our model adopts a hierarchical framework comprising a base model, pattern shifting modules, and consistency enhancement modules, enabling high flexibility and scalability. This design facilitates seamless adaptation to various open-source or proprietary backbone models, such as Flux, Kontext, HiDream, and Qwen-Image. For each DiT-Block, a LoRA layer is inserted at the QKV Attention position as the pattern shifting module, and another two LoRA layers are added at the Projections following Attention as the consistency enhancement module. During training, all original parameters of the backbone model are frozen, and only the parameters of the pattern shifting and consistency enhancement modules are optimized.
@@ 4.2 Quantitative Evaluation
To assess the model’s performance on e-commerce editing tasks, we established an e-commerce evaluation benchmark, EcomEdit-Bench, inspired by widely adopted community standards such as GEdit-Bench. This benchmark encompasses a variety of common e-commerce scenarios, including portrait pose editing, background replacement, accessory addition, hairstyle modification, addition, removal, and replacement of e-commerce elements, text deletion or modification, watermark removal, lighting adjustment, and more, with a total of 380 image-instruction pairs. We employ an automated and objective evaluation protocol to measure the model’s effectiveness. For each generated image, GPT-4o is used to score three aspects: instruction adherence, image quality, and consistency and calculate the average VIE Score .

### v2

@@ 2.2 Editing Pairs Construction Methods
Constructing editing pairs from collected raw image data is the core process in building image editing datasets. Given the wide variety of image editing tasks, each with distinct characteristics and requirements, we systematically categorize these tasks and, accordingly, establish four different methods for constructing editing pairs: the Editing Expert Model Based Method, the Template Based Method, the In-Context Generation Based Method, and the LoRA Hu et al. (2022) Based Method.
@@ 2.2.3 In-Context Generation Based Method
2.2.4 LoRA Based Method
@@ 3.1 Model Architecture
Our model adopts a hierarchical framework comprising a base model, pattern shifting modules, and consistency enhancement modules, enabling high flexibility and scalability. This design facilitates seamless adaptation to various open-source or proprietary backbone models, such as Flux, Kontext, HiDream Cai et al. (2025), and Qwen-Image. For each DiT-Block, a LoRA layer is inserted at the QKV Attention position as the pattern shifting module, and another two LoRA layers are added at the Projections following Attention as the consistency enhancement module. During training, all original parameters of the backbone model are frozen, and only the parameters of the pattern shifting and consistency enhancement modules are optimized.
@@ 4.2 Quantitative Evaluation
To assess the model’s performance on e-commerce editing tasks, we established an e-commerce evaluation benchmark, EcomEdit-Bench, inspired by widely adopted community standards such as GEdit-Bench Liu et al. (2025). This benchmark encompasses a variety of common e-commerce scenarios, including portrait pose editing, background replacement, accessory addition, hairstyle modification, addition, removal, and replacement of e-commerce elements, text deletion or modification, watermark removal, lighting adjustment, and more, with a total of 380 image-instruction pairs. We employ an automated and objective evaluation protocol to measure the model’s effectiveness. For each generated image, GPT-4o is used to score three aspects: instruction adherence, image quality, and consistency and calculate the average VIE Score .